how many moles and molecules of o2 are there in 64g of O2 .What is the mass of one molecules of O2?
1
Expert's answer
2013-08-16T11:12:42-0400
Moles are in O2 in 64 grams:
n = m/M
= 64g/32g per mol
= 2 moles
Molecule of O2 consisy of 2 anoms, then:
64 g x 1 mol/ 16 g = 64 mol / 16 (grams cancel)
64 / 16 = 4 mol
Then in order to find molecules you need to know avogadro's number, which is basically just the definition and value for the unit mol.
avogadro's number states that
1 mole = 6.02 x 10^23
so, we use a conversion factor again to find how many molecules are in 4 moles.
4 moles X 6.02 X 10^23 / 1 mole = (moles cancel)(multiply 4 and 6.02) = 24.08 x 10^23
You need to move the decimal point so that the number is between 1 and 10. So we move the decimal point once space to the left to create: 2.408 The exponent also needs to change from 10^23 to 10^24 to account for the decimal place we moved. So, in 64 grams of oxygen there are 2.408 x 10^24 oxygen molecules, you should probably just write this answer as 2.4 x 10^24 because this accounts for the significant figures you started out with (you had 2 when we started out with the molar mass of oxygen @ 16) and now, we have two again.

Answer: 2.4 x 10^24